OpenStudy (anonymous):

what equation of the line passing trhough the points (-2,3) and (1,4)?

OpenStudy (anonymous):

y=mx+c m=(y2-y1)/(x2-x1) m=4-3/1+2 m=1/3 y=1/3x+c now put any point in this eq and find value of c 4=1/3+c c=11/3 put this value y=1/3x+11/3